What is the energy of photons that corresponds to the wave number of 2.5*10-5 cm-

Grade:12th pass

1 Answers

sonika p
6 years ago
given wave number = 2.5*10-5 cm-1 = 2.5*10-3 m-1
energy of photons is given by E =hc*wave_number
E = 6.636*10^-34 * 3*10^8 * 2.5* 10-3 = 4.969*10^-28 Joules